(1 point) Find the area of the surface obtained by rotating the curve 4x = y2 + 8 about x-axis from x = 2 to x = 4. Area:
Answer #1

4x = y2 +8 phone = 4x–8 y= 14x-8 dy 4 ax 2/4-8 2 √4x-8 Then do = - + The surface area obtained by rotating the curve about x